OpenStudy (anonymous):

(x+5)/(x^2-2x-3)-(x-4)/(x^2-3x-4). Can you show me how to perform the indicated operations step-by-step?

OpenStudy (anonymous):

you can't solve it but you can simplify it

OpenStudy (anonymous):

\[\frac{x+5}{x^2-2x-3} - \frac{x-4}{x^2-3x-4}\] \[\frac{x^2-3x-4}{x^2-3x-4} \cdot \frac{x+5}{x^2-2x-3} - \frac{x-4}{x^2-3x-4} \cdot \frac{x^2-2x-3}{x^2-2x-3}\] \[\frac{(x^2-3x-4)(x + 5) - (x - 4)(x^2-2x-3)}{(x^2-3x-4)(x^2-2x-3)}\] Now try to simplify it.

OpenStudy (anonymous):

Try simplify the entire numerator then factor it and then remove common factors.

OpenStudy (anonymous):

\[\frac{ 8 }{ x^2-2x-3 }\]

OpenStudy (anonymous):

Yes, this is correct.